Grade 12th passMechanics

the vapour density of pcl5 is 104.16 but when heated to 230.c its vapour density is reduced to 62. the degree os dissociation of pcl5 at this temperature will be

Tejaswini
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
as you know , i = 1+ (y-1)α...α = degree of dissociation
y= no of moles of products from one mole reactant
for PCl5 , y = 2
=> i = 1+α
also, y = (normal v.d) /(abnormal v.d.) = 104.16 / 62 = 1.68
= >α = 0.68

Regards
Arun (askIITians forum expert)